OpenStudy (anonymous):

Zach borrowed four thousand dollars from his grandfather to buy a used car so he could travel to and from work. After working for 3 years, he has saved five thousand four hundred forty dollars to pay back to his grandfather. At what interest rate did Zach borrow the money?

OpenStudy (anonymous):

@greenlegodude57

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@greenlegodude57

OpenStudy (anonymous):

Do you know how to set up the problem?

OpenStudy (anonymous):

Borrowed 4000 Worked 3 years Payed back 5440

OpenStudy (anonymous):

You have to find the interest rate in 3 years from 4000 to 5440.

OpenStudy (anonymous):

5440 / 4000 is approximately 1.36 1.36 / 3 is approximately 0.45 So the interest rate is 0.45%. Check: .45 x 3 = 1.35 1.36 * 4000 = 5400. That's close enough.

OpenStudy (anonymous):

Thats it

OpenStudy (anonymous):

There isn't a choice 0.45. :/

OpenStudy (anonymous):

I think it's .25, because .25 * 5440 = 1360 and 5440 - 1360 = 4080.

OpenStudy (anonymous):

Thanks.